OpenStudy (anonymous):

What number increased by 1/5 of itself equals 84?

OpenStudy (anonymous):

So, the equation you need to set up is x+1/5x=84. Solve that. So that would be 6/5x=84, so 70 is the answer. You can check it too. 70+(1/5)*70= 84

OpenStudy (anonymous):

Can someone answer my question here: http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e3967290b8bf47d065fb2bc

OpenStudy (anonymous):

let the number be x then \[x + \left( 1\div5 \right)x = 84\] now LCM \[\left( 5x + x \right)\div5 = 84\] \[6x \div 5 = \] \[6x = 84\times5\] \[6x = 420\] \[x = 420\div6\] \[x = 70\]